# Benchmark of segmentation techniques for pelvic fracture in CT and X-ray: Summary of the PENGWIN 2024 challenge

### v1

@@ Abstract
The segmentation of pelvic fracture fragments in CT and X-ray images is crucial for trauma diagnosis, surgical planning, and intraoperative guidance. However, accurately and efficiently delineating the bone fragments remains a significant challenge due to complex anatomy and imaging limitations. The PENGWIN challenge, organized as a MICCAI 2024 satellite event, aimed to advance automated fracture segmentation by benchmarking state-of-the-art algorithms on these complex tasks. A diverse dataset of 150 CT scans was collected from multiple clinical centers, and a large set of simulated X-ray images was generated using the DeepDRR method. Final submissions from 16 teams worldwide were evaluated under a rigorous multi-metric testing scheme. The top-performing CT algorithm achieved an average fragment-wise intersection over union (IoU) of 0.930, demonstrating satisfactory accuracy. However, in the X-ray task, the best algorithm attained an IoU of 0.774, highlighting the greater challenges posed by overlapping anatomical structures. Beyond the quantitative evaluation, the challenge revealed methodological diversity in algorithm design. Variations in instance representation, such as primary-secondary classification versus boundary-core separation, led to differing segmentation strategies. Despite promising results, the challenge also exposed inherent uncertainties in fragment definition, particularly in cases of incomplete fractures. These findings suggest that interactive segmentation approaches, integrating human decision-making with task-relevant information, may be essential for improving model reliability and clinical applicability.
keywords:
Pelvic fracture, Image segmentation, Deep learning, Challenge
journal: Medical Image Analysis
[…]
[Rossum]organization=Beijing Rossum Robot Technology Co., Ltd., city=Beijing, country=China
[…]
[Beihang]organization=Key Laboratory of Biomechanics and Mechanobiology, Ministry of Education, Beijing Advanced Innovation Center for Biomedical Engineering, School of Biological Science and Medical Engineering, Beihang University, city=Beijing, postcode=100083, country=China
[…]
[JHU]organization=Department of Computer Science, Johns Hopkins University, city=Baltimore, state=MD, country=USA
[…]
[DKFZ1]organization=Division of Medical Image Computing, German Cancer Research Center (DKFZ), city=Heidelberg, country=Germany
[…]
[DKFZ2]organization=Helmholtz Imaging, German Cancer Research Center (DKFZ), city=Heidelberg, country=Germany
[…]
[SMILE]organization=Smart Medical Imaging, Learning and Engineering (SMILE) Lab, Medical UltraSound Image Computing (MUSIC) Lab, School of Biomedical Engineering, Shenzhen University Medical School, Shenzhen University, city=Shenzhen, country=China
[…]
[MedIG]organization=Institute of Biomedical and Health Engineering, Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences, city=Shenzhen, country=China
[…]
[LEM]organization=Pattern Recognition Lab, Friedrich-Alexander-University Erlangen-Nuremberg, city=Erlangen, country=Germany
[…]
[AGH1]organization=Department of Measurement and Electronics, AGH University of Krakow, city=Krakow, country=Poland
[…]
[AGH2]organization=MedApp S.A., city=Krakow, country=Poland
[…]
[XMU]organization=National Institute for Data Science in Health and Medicine, Xiamen University, city=Xiamen, country=China
[…]
[Sano1]organization=Faculty of Informatics, Mathematics, and Mechanics, University of Warsaw, city=Warsaw, country=Poland
[…]
[Sano2]organization=Faculty of Electronics and Information Technology, Warsaw University of Technology, city=Warsaw, country=Poland
[…]
[Sano3]organization=Sano Center for Computational Medicine, city=Krakow, country=Poland
[…]
[UCLA]organization=Department of Radiation Oncology, University of California, Los Angeles, city=Los Angeles, state=CA, country=USA
[…]
[USTC]organization=School of Biomedical Engineering, Division of Life Sciences and Medicine, University of Science and Technology of China, city=Hefei, country=China
[…]
[JST]organization=Department of Orthopaedics and Traumatology, Beijing Jishuitan Hospital, city=Beijing, country=China
supportedsupportedfootnotetext: This work was supported by Nation Natural Science Foundation of China (NSFC6247010104), National Key Research and Development Program of China (2022YFC2504304), Natural Science Foundation of Beijing (L222136), and Beijing Science and Technology Project (Z221100003522007 and Z241100009024030)
1 Introduction
@@
Pelvic fractures are among the most severe orthopedic injuries, often resulting from high-energy trauma such as motor vehicle accidents and falls from height. These fractures are associated with a disability rate exceeding 50% and a mortality rate above 13%, making them the deadliest of all compound fractures (Kowal et al., 2007; Sugano, 2013). The complex anatomy of the pelvis, combined with surrounding soft tissues, makes surgical intervention particularly challenging. In recent years, robotic-assisted closed fracture reduction surgeries have demonstrated improved clinical outcomes with reduced bleeding, lower infection rates, decreased vascular and nerve damage, and faster postoperative recovery (Ge et al., 2022). The precision and outcome of these image-guided robotic surgery heavily rely on accurate preoperative and intraoperative imaging and segmentation.
@@
Automated pelvic fracture segmentation is a crucial step in trauma diagnosis and surgical planning. In 3D CT scans, precise fracture segmentation is a prerequisite for fragment-wise reduction planning, enabling accurate pose and trajectory planning, as well as screw planning, for surgical realignment and fixation (Liu et al., 2025; Shi et al., 2025). It also plays a key role in and supports detailed fracture classification, aiding in clinical decision-making. In 2D C-arm X-ray imaging, fracture segmentation is equally important for diagnostic interpretation and intra-operative guidance. By segmenting fractures in radiological images, surgical plans can be accurately transferred to the operating space, facilitating C-arm imaging-based navigation and registration (Han et al., 2021).
@@
Traditionally, fracture segmentation in CT has been performed semi-automatically using 3D interactive tools. In the current surgical planning workflow for robotic pelvic fracture reduction, the fractured pelvis is typically extracted using a graph-cut algorithm, with manually selected foreground and background seed points. Following this step, each bone fragment must be individually isolated. For completely separated fragments, connected component analysis can be used for extraction. However, when fragments remain connected or collided, operators must carefully delineate the fracture surfaces, either by manually tracing contours slice by slice in 2D images or by examining the fracture structure in a 3D view. This process is highly labor-intensive, often requiring 0.5 to 1 hour for complex fractures with overlapping surfaces, not to mention the additional time required for subsequent fragment pose planning. The efficiency of this workflow is particularly critical, as pelvic fracture reductions are typically non-elective procedures, sometimes required in emergency settings. Moreover, the segmentation accuracy is heavily dependent on the skill and experience of the operator.
@@
For intraoperative X-ray imaging, fracture segmentation holds significant potential to facilitate pelvic reduction surgery; however, it has not yet been adopted into clinical routine. Pelvic fracture fixation is a highly complex procedure that relies on radiographic imaging to accurately place Kirschner wires and orthopedic screws within specific bony corridors, following a detailed preoperative plan. In intraoperative cone-beam CT (CBCT) imaging, efforts have been successfully made to automatically extract bone regions and register them with preoperative CT scans (Liu et al., 2024). However, similar techniques for X-ray imaging remain underexplored. While C-arm X-ray imaging is far more widely used than CBCT, its segmentation process presents several challenges, including the difficulty in distinguishing overlapping pelvic structures, the limited visibility of fracture lines, and the need for precise alignment of fractured fragments. These challenges highlight the need for improved segmentation techniques in X-ray imaging to enhance intraoperative guidance with lower radiation exposure and wider applicability.
@@
The Pelvic Bone Fragments with Injuries (PENGWIN) Challenge 2024 was launched to address these challenges by fostering the development of robust, automated segmentation techniques for pelvic fractures in both 3D CT and 2D X-ray modalities. Hosted as a satellite event at the 27th International Conference on Medical Image Computing and Computer Assisted Intervention (MICCAI 2024), this challenge aimed to:
  1. 1.
    Advance the state of the art in pelvic fracture segmentation using deep learning.
  2. 2.
    Provide a high-quality, publicly available dataset to facilitate further research in medical image analysis and orthopedics.
  3. 3.
    Promote interdisciplinary collaboration between researchers in medical imaging, artificial intelligence, and orthopedic surgery.
To achieve these goals, we curated a comprehensive, multi-institutional dataset consisting of CT scans from 150 patients scheduled for pelvic reduction surgery. These scans were acquired from multiple research centers using various scanning devices to ensure diversity in imaging conditions and patient cohorts. The ground-truth segmentations of the hipbones, sacrum, and fracture fragments were semi-automatically annotated and validated by medical experts. In addition, we generated 48,600 high-quality, realistic X-ray images using the DeepDRR method (Unberath et al., 2018), simulating diverse virtual C-arm angles and incorporating surgical tool interference to closely reflect real-world intraoperative conditions. Participants were challenged to develop fully automated segmentation algorithms for 3D CT and 2D X-ray tasks, contributing novel methods that improve the accuracy, robustness, and clinical applicability of fracture segmentation.
@@
2 Related works
2.1 Orthopedic image processing challenge
@@
Table 1: Summary of major orthopedic image processing algorithmic challenges between 2017 and 2024. Challenges directly related to bone fractures are shown in bold.
Acronym Challenge title Organization Year Modality Anatomy Task Data size
PBA-RSNA (Halabi et al., 2019) RSNA pediatric bone age challenge RSNA 2017 X-ray Hand Scoring 14,236
AASCE (Wang et al., 2021) Accurate automated spinal curvature estimation MICCAI 2019 X-ray Spine Measurement 707
VerSe’19 (Sekuboyina et al., 2021) Large scale vertebrae segmentation challenge 2019 MICCAI 2019 CT Spine Identification, segmentation 160
VerSe’20 (Sekuboyina et al., 2021) Large scale vertebrae segmentation challenge 2020 MICCAI 2020 CT Spine Segmentation 374
RA2-DREAM (Sun et al., 2022) Automated scoring of radiographic joint damage DREAM 2020 X-ray Hand, foot Scoring 698
object-CXR Automatic detection of foreign objects on chest X-rays MIDL 2020 X-ray Chest Identification 10,000
KNOAP2020 (Hirvasniemi et al., 2023) Knee osteoarthritis prediction challenge - 2020 MRI, X-ray Knee Classification 423
RibFrac (Yang et al., 2024) Rib fracture detection and classification challenge MICCAI 2020 CT Rib Identification, classification 660
AutoImplant (Li et al., 2021) Automatic cranial implant design challenge MICCAI 2020 Binary volume Cranium Implant design 200
AutoImplant2021 (Li et al., 2023) Automatic cranial implant design challenge 2021 MICCAI 2021 Surface model Cranium Implant design 681
CSF-RSNA (Hu et al., 2024) RSNA cervical spine fracture AI challenge RSNA 2022 CT Cervical spine Identification 1,400
K2S (Tolpadi et al., 2023) From undersampled k-space to automatic segmentation MICCAI 2022 MRI Knee Segmentation 350
FBD-Kaggle (Khanal et al., 2023) Fractured bone detection challenge - 2023 CT Limb Identification 24
CL-Detection (Wang et al., 2016) Cephalometric landmark detection in lateral X-ray Images MICCAI 2023 X-ray Skull Identification 600
LSDC-RSNA (Richards et al., 2024) RSNA lumbar spine degenerative classification AI challenge RSNA 2024 MRI Lumbar spine Classification 1,975
SPIDER (van der Graaf et al., 2024) Spine segmentation: discs, vertebrae, and spinal canal - 2024 MRI Lumbar spine Segmentation 544
We review the algorithmic challenges related to orthopedic image processing that took place between 2017 and 2024. Our search scope includes: (1) challenges endorsed by major medical imaging conferences, such as MICCAI, the IEEE International Symposium on Biomedical Imaging (ISBI), and Medical Imaging with Deep Learning (MIDL); and (2) publicly accessible challenges from key online platforms, including the Grand Challenge platform, Dialogue for Reverse Engineering Assessments and Methods (DREAM) Challenges, Kaggle Competitions, and Radiological Society of North America (RSNA) Challenges. This list of registries was compiled with reference to the MICCAI Special Interest Group (SIG) for Challenges. The 16 challenges surveyed are summarized in Table 1. Furthermore, challenges that are directly related to bone fractures are reviewed in detail.
@@
The RibFrac Challenge (Yang et al., 2024) was launched with MICCAI 2020, aimed to advance automated rib fracture detection and diagnosis. The dataset comprises over 5,000 rib fractures across 660 CT scans. The challenge included two tasks. In the first task, fracture detection was formulated as an instance segmentation problem, with annotated fractured regions provided as ground truth. Due to the inherent ambiguity in delineating fracture boundaries, the segmentation outputs were primarily used to evaluate detection performance, with the free-response receiver operating characteristic (FROC) analysis employed for assessment. In the second task, the detected rib fractures were further classified into four clinically relevant categories: buckle fractures, non-displaced fractures, displaced fractures, and segmental fractures. Classification performance was evaluated using the F1 score.
@@
The 2022 RSNA Cervical Spine Fracture AI Challenge (Hu et al., 2024) focused on rapid detection and localization of cervical spine fractures, which is critical to prevent neurological deterioration and paralysis following trauma. The dataset includes over 3,000 CT scans, with an approximately equal distribution of positive and negative cases. Participants were tasked with predicting the presence or absence of fractures at each of the seven cervical vertebrae, but were not required to locate them within the CT volume. Submissions were evaluated based on a weighted multi-label logarithmic loss, assessing the predicted probability of fracture at each vertebra.
@@
The Fractured Bone Detection Challenge (Khanal et al., 2023) was a lightweight, independent challenge launched on Kaggle. It provided a small dataset of 24 CT scans covering both upper and lower limbs. The objective was to identify fractures in 2D CT slices to facilitate rapid diagnosis. Each CT slice was binary classified based on the presence of fractures, with area under the sensitivity-specificity curve (AUC) used as the evaluation metric.
@@
In summary, challenges focusing on orthopedic fractures remain limited in number, with no prior challenge specifically targeting pelvic bone or X-ray segmentation. Existing efforts have predominantly focused fracture detection and localization at a coarse level. However, the emerging adoption of image-guided reduction surgery highlights the growing demand for voxel-wise fracture fragment segmentation, calling for new challenges and benchmarks to drive technical advancements in this domain.
2.2 Automated fracture segmentation
2.2.1 Fracture segmentation in CT
@@
Traditional non-learning-based methods for automated and semi-automated CT fracture segmentation have been comprehensively summarized in previous studies (Liu et al., 2025). These approaches include intensity thresholding, graph cut, and region growing methods. While such methods can be effective in cases involving non-contact fractures, they often struggle in more complex scenarios, such as fractures with fragment contact, impaction, or compression (Krčah et al., 2011). In contrast, deep learning-based fracture segmentation remains a relatively underexplored area; however, recent studies have demonstrated its significant potential. In the following, we review the latest advancements in learning-based approaches for automated fracture segmentation in CT.
@@
Early studies have explored the direct application of established segmentation networks to fracture fragment segmentation in CT scans. Yang et al. (2022) applied a Mask R-CNN-based instance segmentation model to locate and extract individual fragments from 2D CT slices of intertrochanteric fractures. Kim et al. (2023) utilized DeepLab v3+ for semantic segmentation of tibial and fibular fracture fragments in 2D CT slices. In their approach, fragment labels were manually assigned during data preparation; however, this strategy raises concerns regarding robustness due to potential label inconsistencies across patients and the limited scalability of treating individual fragments as semantic classes. Wang et al. (2022a) adopted a 3D V-Net model to segment intertrochanteric femoral fractures, using a three-class labeling scheme to distinguish the proximal femur, fracture fragment, and distal femur.
@@
In our previous study, FracSegNet, fracture segmentation is performed using a two-stage pipeline comprising two sequential networks (Liu et al., 2023). First, an anatomical segmentation network extracts the hipbones and sacrum from CT images. Subsequently, a fracture segmentation network differentiates the primary fragments (label 1) and secondary fragments (label 2) within each bone region. The fracture segmentation network is trained using a distance-weighted loss, which emphasizes attention near fracture boundaries and improves performance in these critical regions. This approach enables semantic segmentation of both primary and secondary fragments, thereby avoiding the need to explicitly handle a variable number of fragments during network inference. Following segmentation, secondary fragments are further separated using connected component analysis. However, this post-processing step is insufficient when secondary fragments are in contact, leading to failures in accurately separating adjoining fragments.
@@
Building upon the same primary-secondary representation, Zeng et al. (2024) proposed a pelvic fracture segmentation network that incorporated the prediction of a fragment prior distance map as an auxiliary task to better capture the spatial relationships between fragments. In addition, a dynamic multi-scale feature fusion module was introduced to address the challenge posed by the varying sizes of fragments.
@@
In addition, Deng et al. (2023) explored the synergistic integration of fragment segmentation and reduction pose planning for femoral fractures. Their method takes as input segmented bone surface point clouds from both the fractured and contralateral healthy sides. By using the healthy side as a template, the segmentation of fracture fragments is refined through an alternating process that minimizes the distance between the current fracture model and the template. However, this approach is only applicable when a complete contralateral template is available and may be limited by natural anatomical asymmetry.
2.2.2 Fracture segmentation in X-ray

### v2

@@ Abstract
The segmentation of pelvic fracture fragments in CT and X-ray images is crucial for trauma diagnosis, surgical planning, and intraoperative guidance. However, accurately and efficiently delineating the bone fragments remains a significant challenge due to complex anatomy and imaging limitations. The PENGWIN challenge, organized as a MICCAI 2024 satellite event, aimed to advance automated fracture segmentation by benchmarking state-of-the-art algorithms on these complex tasks. A diverse dataset of 150 CT scans was collected from multiple clinical centers, and a large set of simulated X-ray images was generated using the DeepDRR method. Final submissions from 16 teams worldwide were evaluated under a rigorous multi-metric testing scheme. The top-performing CT algorithm achieved an average fragment-wise intersection over union (IoU) of 0.930, demonstrating satisfactory accuracy. However, in the X-ray task, the best algorithm achieved an IoU of 0.774, which is promising but not yet sufficient for intra-operative decision-making, reflecting the inherent challenges of fragment overlap in projection imaging. Beyond the quantitative evaluation, the challenge revealed methodological diversity in algorithm design. Variations in instance representation, such as primary-secondary classification versus boundary-core separation, led to differing segmentation strategies. Despite promising results, the challenge also exposed inherent uncertainties in fragment definition, particularly in cases of incomplete fractures. These findings suggest that interactive segmentation approaches, integrating human decision-making with task-relevant information, may be essential for improving model reliability and clinical applicability.
I Introduction
@@
Pelvic fractures are among the most severe orthopedic injuries, often resulting from high-energy trauma such as motor vehicle accidents and falls from height. These fractures are associated with a disability rate exceeding 50% and a mortality rate above 13%, making them the deadliest of all compound fractures [kowal2007basics, sugano2013computer]. The complex anatomy of the pelvis, combined with surrounding soft tissues, makes surgical intervention particularly challenging. In recent years, robotic-assisted closed fracture reduction surgeries have demonstrated improved clinical outcomes with reduced bleeding, lower infection rates, decreased vascular and nerve damage, and faster postoperative recovery [LSBT2022robot]. The precision and outcome of these image-guided robotic surgery heavily rely on accurate preoperative and intraoperative imaging and segmentation.
@@
Automated pelvic fracture segmentation is a crucial step in trauma diagnosis and surgical planning. In 3D CT scans, precise fracture segmentation is a prerequisite for fragment-wise reduction planning, enabling accurate pose and trajectory planning, as well as screw planning, for surgical realignment and fixation [liu2025preoperative, shi2025automatic]. It also plays a key role in detailed fracture classification, aiding in clinical decision-making. In 2D C-arm X-ray imaging, fracture segmentation is equally important for diagnostic interpretation and intra-operative guidance. By segmenting fractures in radiological images, surgical plans can be accurately transferred to the operating space, facilitating C-arm imaging-based navigation and registration [han2021fracture].
@@
Traditionally, fracture segmentation in CT has been performed semi-automatically using 3D interactive tools. In the current surgical planning workflow for robotic pelvic fracture reduction, the fractured pelvis is typically extracted using a graph-cut algorithm, with manually selected foreground and background seed points. Following this step, each bone fragment must be individually isolated. For completely separated fragments, connected component analysis can be used for extraction. However, when fragments remain connected or collided, operators must carefully delineate the fracture surfaces, either by manually tracing contours slice by slice in 2D images or by examining the fracture structure in a 3D view. This process is highly labor-intensive, often requiring 0.5 to 1 hour for complex fractures with overlapping surfaces, not to mention the additional time required for subsequent fragment pose planning. The efficiency of this workflow is particularly critical, as pelvic fracture reductions are often non-elective procedures, sometimes required in emergency settings. Moreover, the segmentation accuracy is heavily dependent on the skill and experience of the operator.
@@
For intraoperative X-ray imaging, fracture segmentation holds significant potential to facilitate pelvic reduction surgery; however, it has not yet been adopted into clinical routine. Pelvic fracture fixation is a highly complex procedure that relies on radiographic imaging to accurately place Kirschner wires and screws within specific bony corridors, following a detailed preoperative plan. In intraoperative cone-beam CT (CBCT) imaging, efforts have been successfully made to automatically extract bone regions and register them with preoperative CT scans [liu2024automatic]. However, similar techniques for X-ray remain underexplored. While C-arm X-ray imaging is far more widely used than CBCT, its segmentation process presents several challenges, including the difficulty in distinguishing overlapping structures, the limited visibility of fracture lines, and the need for precise alignment of fractured fragments. These challenges highlight the need for improved segmentation techniques in X-ray to enhance intraoperative guidance with lower radiation exposure and wider applicability.
@@
The Pelvic Bone Fragments with Injuries (PENGWIN) Challenge 2024 was launched to address these challenges by fostering the development of robust, automated segmentation techniques for pelvic fractures in both 3D CT and 2D X-ray modalities. Hosted as a satellite event at the 27th International Conference on Medical Image Computing and Computer Assisted Intervention (MICCAI 2024), this challenge aimed to: (1) advance the state of the art in pelvic fracture segmentation using deep learning; (2) provide a high-quality, publicly available dataset to facilitate further research in medical image analysis and orthopedics; and (3) promote interdisciplinary collaboration between researchers in medical imaging, artificial intelligence, and orthopedic surgery.
To achieve these goals, we curated a comprehensive, multi-institutional dataset consisting of CT scans from 150 patients scheduled for pelvic reduction surgery. These scans were acquired from multiple research centers using various scanning devices to ensure diversity in imaging conditions and patient cohorts. The ground-truth segmentations of the hipbones, sacrum, and fracture fragments were semi-automatically annotated and validated by medical experts. In addition, we generated 48,600 high-quality, realistic X-ray images using the DeepDRR method [unberath2018deepdrr] , simulating diverse virtual C-arm angles and incorporating surgical tool interference to closely reflect real-world intraoperative conditions. Participants were challenged to develop fully automated segmentation algorithms for 3D CT and 2D X-ray tasks, contributing novel methods that improve the accuracy, robustness, and clinical applicability of fracture segmentation.
@@
II Related Works
II-A Challenges on Orthopedic Fracture
@@
We surveyed algorithmic challenges related to orthopedic image processing between 2017 and 2024. Challenges directly related to bone fractures are reviewed in detail.
@@
The RibFrac Challenge [yang2024deep] was launched with MICCAI 2020, aimed to advance automated rib fracture detection and diagnosis. The dataset comprises over 5,000 rib fractures across 660 CT scans. The challenge included two tasks. In the first task, fracture detection was formulated as an instance segmentation problem, with annotated fractured regions provided as ground truth. Due to the inherent ambiguity in delineating fracture boundaries, the segmentation outputs were primarily used to evaluate detection performance, with FROC analysis employed for assessment. In the second task, the detected rib fractures were further classified into four clinically categories: buckle fractures, non-displaced fractures, displaced fractures, and segmental fractures.
@@
The 2022 RSNA Cervical Spine Fracture AI Challenge [hu2024assessing] focused on rapid detection and localization of cervical spine fractures, which is critical to prevent neurological deterioration and paralysis following trauma. The dataset includes over 3,000 CT scans, with an approximately equal distribution of positive and negative cases. Participants were tasked with predicting the presence or absence of fractures at each of the seven cervical vertebrae, but were not required to locate them within the CT volume.
@@
The Fractured Bone Detection Challenge [fractured-bone-detection-challenge] was a lightweight, independent challenge launched on Kaggle. It provided a small dataset of 24 CT scans covering both upper and lower limbs. The objective was to identify fractures in 2D CT slices to facilitate rapid diagnosis. Each CT slice was binary classified based on the presence of fractures, with the area under curve (AUC) used as the evaluation metric.
@@
In summary, challenges focusing on fractures remain limited in number, with no prior challenge specifically targeting pelvic bone or X-ray segmentation. Existing efforts have predominantly focused on fracture detection at a coarse level. However, the emerging adoption of image-guided reduction surgery highlights the growing demand for voxel-wise fracture fragment segmentation, calling for new challenges and benchmarks to drive advancements in this domain.
II-B Automated Fracture Segmentation
II-B1 Fracture Segmentation in CT
@@
Traditional non-learning-based methods for automated and semi-automated CT fracture segmentation have been comprehensively summarized in previous studies [liu2025preoperative]. These approaches include intensity thresholding, graph cut, and region growing methods. While such methods can be effective in cases involving non-contact fractures, they often struggle in more complex scenarios, such as fractures with fragment contact, impaction, or compression [2011fullyMaxFlow]. In contrast, deep learning-based fracture segmentation remains a relatively underexplored area; however, recent studies have demonstrated its significant potential. In the following, we review the latest advancements in learning-based approaches for automated fracture segmentation in CT.
@@
Early studies have explored the direct application of established segmentation networks to fracture fragment segmentation in CT scans. Yang et al. [yang2022recognition] applied a Mask R-CNN-based instance segmentation model to locate and extract individual fragments from 2D CT slices of intertrochanteric fractures. Kim et al. [kim2023automatic] utilized DeepLab v3+ for semantic segmentation of tibial and fibular fragments in 2D CT slices. In their approach, fragment labels were manually assigned during data preparation; however, this strategy raises concerns regarding robustness due to potential label inconsistencies across patients and the limited scalability of treating individual fragments as semantic classes. Wang et al. [wang2022accuracy] adopted a 3D V-Net model to segment intertrochanteric femoral fractures, using a three-class labeling scheme to distinguish the proximal femur, fracture fragment, and distal femur.
@@
In our previous study, FracSegNet, fracture segmentation is performed using a two-stage pipeline comprising two sequential networks [liu2023pelvic]. First, an anatomical segmentation network extracts the hipbones and sacrum from CT images. Subsequently, a fracture segmentation network differentiates the primary fragment and secondary fragments within each bone region. The fracture segmentation network is trained using a distance-weighted loss, which emphasizes attention near fracture boundaries and improves performance in these critical regions. This approach enables semantic segmentation of major and minor fragments, thereby avoiding the need to explicitly handle a variable number of fragments during network inference. Following segmentation, secondary fragments are further separated using connected component analysis. However, this post-processing step can be insufficient when secondary fragments are in contact, leading to failures in accurately separating small fragments.
@@
Building upon the same primary-secondary representation, Zeng et al. [zeng2024fragment] proposed a pelvic fracture segmentation network that incorporated the prediction of a fragment prior distance map as an auxiliary task to better capture the spatial relationships between fragments. In addition, a dynamic multi-scale feature fusion module was introduced to address the challenge posed by the varying sizes of fragments.
@@
In addition, Deng et al. [deng2023synergistically] explored the synergistic integration of fragment segmentation and reduction pose planning for femoral fractures. Their method takes as input segmented bone surface point clouds from both the fractured and contralateral healthy sides. By using the healthy side as a template, the segmentation of fracture fragments is refined through an alternating process that minimizes the distance between the current fracture model and the template. However, this approach is only applicable when a complete contralateral template is available and may be limited by natural anatomical asymmetry.
II-B2 Fracture Segmentation in X-Ray